OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, -1) and perpendicular to y = 1/2x + 3. y = -2x - 7 y = -2x +5 y = 2x + 5 y = 2x - 7

OpenStudy (anonymous):

slope is negative reciprocal m = -2 -1 = -2(-3) +b -7 = b y = -2x - 7 y = -(2x + 7)

OpenStudy (anonymous):

So it would be the last option cause of the negative sign in front of the parenthesis ?

OpenStudy (anonymous):

@mattt9

OpenStudy (anonymous):

no it's the first option. The way of writing it with the parentheses is just the simplified form. The negative sign is distributed to both terms within the parentheses

OpenStudy (anonymous):

Thank you